OpenStudy (anonymous):

55 mi / h = ____ ft / s

OpenStudy (help!!!!):

55*5280=290,400ft/s

OpenStudy (mathstudent55):

@HELP!!!! That is ft/h. The question is ft/s. \(55 \dfrac{mi}{h} \times \dfrac{5280~ft}{1~mi} \times \dfrac{1~h}{60~mn} \times \dfrac{1~mn}{60~s} \) Let's check the units: \(55 \dfrac{\cancel{\color{red}{mi}}}{\cancel{\color{blue}{h}}} \times \dfrac{5280~ft}{1~\cancel{\color{red}{mi}}} \times \dfrac{1~\cancel{\color{blue}{h}}}{60~\cancel{\color{green}{mn}}} \times \dfrac{1~\cancel{\color{green}{mn}}}{60~s} \) As you can see, all units except ft/s cancel out. Now do the operations.